When a hydrocarbon is burnt completely the ratio of masses of CO2 and H2O formed is 44:27 the hydrocarbon is?

Grade:11

1 Answers

Devang
5 years ago
let assume that   CxH is hydrocarbon
         then
                     CxH  + ( x+y/4) O2  \rightarrow x CO +  y/2 H2O
1 mol of  CO give  44 gram of  CO then
x mol of  CO give  44x  gram of  CO           = 44x  gm  …......(1)
 
1   mol   of   H2O give  18 gram  of  H2O   then
y/2  mol  of  H2O  give  y/2(18)  gram  of  H2O = 9y  gm  ….......(2)
 
we gave ratio of (1) & (2)  is 44:27
then
         44x / 9y = 44 / 27
             x / y   = 1 / 3
then hydrocarbon is  (CH3)n